You are measuring the transport of a drug through an organ via diffusion.

You introduce the drug at a single point in the the tissue using a needle.

If after 15 s the average distance the drug has diffused is 0.51 cm from the initial location how many minutes will it take for the average distance the drug has diffused to be 5.0 cm from the initial location? (to 2 s.f)
1
Expert's answer
2020-04-16T08:45:11-0400
"t=\frac{x^2}{2D }"

Thus,


"t_2=\frac{x^2_2}{x_1^2}t_1=\frac{5^2}{0.51^2}16=1587\ s=26\ min"